OpenStudy (anonymous):

if a differential equations solution divides by x^2 and the initial condition is a point where x=0, what do you do? (problem: dy/dx=xyy or dy/dx=xy^2)

OpenStudy (anonymous):

initial condition is (0,-2)

OpenStudy (anonymous):

You can use separation by variables here and then integrate. Once you have y after integration, you can use the initial condition.

OpenStudy (anonymous):

i did that, but my solution is y=(-2/x^2)+2 which doesnt technically do through the initial condition

OpenStudy (anonymous):

i think its a vertical asymptote which... definitely isint the point XD

OpenStudy (anonymous):

There's nothing wrong with that equation. After you put it into separable form and solve, you should get,\[\frac{x^2y}{2}+cy+1=0\]Initial condition,\[0-2c+1=0 \rightarrow c=\frac{1}{2}\]

OpenStudy (anonymous):

You can then factor out y and rearrange into the form y=y(x).

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

i got c=2?

OpenStudy (anonymous):

When you separated the variable and tried to get y by itself did you include c and also multiply the y by it?

OpenStudy (anonymous):

.... -.-

OpenStudy (anonymous):

ima try that, how about: y'=1+yy

OpenStudy (anonymous):

You can do that problem the same way. I look at it as y' is dy/dx and then separate variables. Make sure when you do though, that the 1 goes with the y^2. So move the entire quantity.

OpenStudy (anonymous):

dang, the small things get by me... that should be something like arctany=x+c? so tan(x+c)=y?

OpenStudy (anonymous):

that explains why my condition is in radians

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

alright... major confusion... \[y'=(e ^{x-y})\div(1+e ^{x})\]

OpenStudy (anonymous):

You want to solve that?

OpenStudy (anonymous):

sadly ><

OpenStudy (anonymous):

Just pull out the e^(-y) and it becomes separable ;)

OpenStudy (anonymous):

asoghaqoeraifoiafauirfhlanihusdfvnnfldanufd

OpenStudy (anonymous):

\[e^ydy=\frac{e^x}{1+e^x}dx\]

OpenStudy (anonymous):

differential equations are gay -.-

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

dont tell me any more!!!

OpenStudy (anonymous):

tell you what? answers?

OpenStudy (anonymous):

yeah.. dont tell me any more lol

OpenStudy (anonymous):

no worries :D

OpenStudy (anonymous):

I think you have the best icon...if that helps...

OpenStudy (anonymous):

hahahaha it was annoyingly hard to make XD

OpenStudy (anonymous):

i have a pretty good feeling that your icon is not you... though there is a significant chance that i am wrong given by the circumstances

OpenStudy (anonymous):

it could be me, since i look like that

OpenStudy (anonymous):

lol, how did your frigging doughnut go?

OpenStudy (anonymous):

pretty good, i have a way of doing things when i need help: use the first sentence or line of the help, fail a thousand times before i use the second, so on and so forth. so i used quite a bit, but not all of your help in the process of doing the assignment, and in the end i was flowing through it quite nicely, i compared my results to yours and they were identical, i then compared my results to the assignment and apparently the last couple steps, which i had done by hand, i could take the integral with a calculator... and then later he gave us the hand written answer, so i felt proud that i did it without a calculator... but yeah your help is insanely apreciated

OpenStudy (anonymous):

alright im stuck at e^y=ln(1+e^x)+c initial condition is (0,1)

OpenStudy (anonymous):

Is that y supposed to be y' (y-prime)?

OpenStudy (anonymous):

nope, thats my solution

OpenStudy (anonymous):

to the above equation

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

Yeah. You're right. Leave it like that too.

OpenStudy (anonymous):

but i need to find c

OpenStudy (anonymous):

Oh, just plug your numbers in.

OpenStudy (anonymous):

\[e=\ln 2 +c \rightarrow c = e- \ln 2\]

OpenStudy (anonymous):

any way to do that without a calc?

OpenStudy (anonymous):

You mean simplify?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

What's wrong with Euler's number? :p

OpenStudy (anonymous):

ah god its 1... can you check these real quick?

OpenStudy (anonymous):

Check what?

OpenStudy (anonymous):

Do you ever sleep?

OpenStudy (anonymous):

not too much anymore lol

OpenStudy (anonymous):

I need paper...I've run out...hang on...

OpenStudy (anonymous):

but all my work is shown?

OpenStudy (anonymous):

I get something different for the first one. Just go to bed and I'll look at it and scan and you can look at it when you get up.

OpenStudy (anonymous):

do i have to multiply the c by 2 aswell?

OpenStudy (anonymous):

I'll do it here:

OpenStudy (anonymous):

my answer doesnt make sense anyways... when i find y' its all weird and quotienty

OpenStudy (anonymous):

\[\frac{dy}{dx}=xy^2 \rightarrow y^{-2}dy = x dx \rightarrow -y^{-1}=x+c\]

OpenStudy (anonymous):

Your boundary condition is (0,-2), so

OpenStudy (anonymous):

x+c or (1/2)xx+c?

OpenStudy (anonymous):

\[-2^{-1}=0+c \rightarrow c=-\frac{1}{2}\]

OpenStudy (anonymous):

So\[y^{-1}+x=\frac{1}{2} \rightarrow 2+2xy=y \rightarrow y(1-2x)=2 \rightarrow y=\frac{2}{1-2x}\]

OpenStudy (anonymous):

waitttt the integral of xdx is (1/2)xx+c though!

OpenStudy (anonymous):

Oh crap, I forgot the frigging x. Sorry Arman.

OpenStudy (anonymous):

nah dont be sorry, i should hit the sheets tho, i'll probably have more questions tomorrow lol

OpenStudy (anonymous):

I still get a different answer to you.

OpenStudy (anonymous):

basic differential equations (like these) arent usually part of calc II are they?

OpenStudy (anonymous):

Yours is close. I'll work through them - they won't take long - and scan and you can read it in the morning.

OpenStudy (anonymous):

different answer?

OpenStudy (anonymous):

I'm not sure about the American system.

OpenStudy (anonymous):

y=2/(1-x^2)

OpenStudy (anonymous):

did you use (0,-2) for the initial whatever?

OpenStudy (anonymous):

pellet, I wrote it wrong on my paper. I wrote (0,2). SORRY AGAIN :'(

OpenStudy (anonymous):

NOW I get what you got.

OpenStudy (anonymous):

>< nah man i hate when ppl tell me sorry

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

my calc teacher, this morning, was like "oh flutter..." *covers mouth*... *whole class starts laughing caus we hear him say pellet sometimes but never the f word so it meant he was really screwed over XD

OpenStudy (anonymous):

lol, it's great when that happens. I got the same results as you for the other two. The *only* thing I would do differently is leave \[e-\log 2\]as is, and not use a numerical approximation (in the last one).

OpenStudy (anonymous):

thats probably better in the math world, where you need to be precise

OpenStudy (anonymous):

Yeah, exactly.

OpenStudy (anonymous):

is that prefered in college?

OpenStudy (anonymous):

preferred*

OpenStudy (anonymous):

Absolutely. You'd only use a numerical approximation for the time you actually need to extract a number. Other than that, we leave it because we may want to do further mathematics on it and we lose information/exactness if we start approximating beforehand.

OpenStudy (anonymous):

just like the AP exams lol. i think i asked you before but what are the remaining major math courses after differential eqas and linear algebra (when i say top i mean like... there should be less than 3 "top" ones)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yeah, I really don't know. Your system is different to the system in UK, Australia and Canada. We don't break mathematics up into sections like calculus and then apply a skill level to it. We apply a skill level to ALL of mathematics at particular points and you learn EVERYTHING at a particular level.

OpenStudy (anonymous):

oh... well what are the chances that i can attain a college grad level (in math) of mathematical understanding in the next year?

OpenStudy (anonymous):

given the rate that i covered calc 1 and 2 within 1.5 months

OpenStudy (anonymous):

I'd say you're pretty good.

OpenStudy (anonymous):

College grad level in calculus...is that what you're asking?

OpenStudy (anonymous):

You have to cover calculus of a complex variable as well. There's also tensor calculus.

OpenStudy (anonymous):

not just calculus, but all mathematics, can i attain that sort of knowledge, about 4 or more years of college math in the next 9 months?

OpenStudy (anonymous):

complex variable wont be too hard, what is tensor calc??

OpenStudy (anonymous):

It *might* take a little longer ;p

OpenStudy (anonymous):

They're devices we use to extend scalars, vectors and matrices to higher dimensions.

OpenStudy (anonymous):

They're very useful.